1 + 2 +4 + 5 +6 +8 ....59 , the sum of series is ? Its answer is 2000
Answer by KMST(5328)

Something is wrong with your problem.
1 + 2 + 3 + 4 + 5 + 6 + 7 + 8 + .... + 59 = 1770.
If you skip 3, 7, 12, 18, and so on the sum will be even less than 1770,
and a lot less than 2000.
In fact, skipping like that will never yield a sum of 2000.
Neither will adding consecutive numbers starting from 1, without skipping any:
1 + 2 + 3 + 4 + 5 + 6 + 7 + 8 + .... + 62 = 1953 and
1 + 2 + 3 + 4 + 5 + 6 + 7 + 8 + .... + 62 + 63 = 2016.